OpenStudy (anonymous):

Would you make more interest by investing $2,500 in a 6-month certificate of deposit (CD) at 3% interest or by investing 1,000 in a 12-Month CD at 4% interest? Show your work

OpenStudy (anonymous):

2500 * 6 * (.03/12)= 1000 *12 *(.04/12 )=1000*1*(.04)=

OpenStudy (anonymous):

1000$ in 12-month checking deoisut at 4 oercent ratio?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

no bank would give you 4% of your interest worth iof 6 month period, what they do is cut the interest ratio in half. so your left with 1.5%

OpenStudy (anonymous):

are you talking about first or second equation